Thanks in advance for helping me. Write the quadratic equation for the set of roots 4, -6
Answer by ankor@dixie-net.com(22740)

Write the quadratic equation for the set of roots 4, -6
:
x = 4
(x-4) = 0
and
x = -6
(x+6) = 0
:
FOIL: (x+6)(x-4) = x^2 - 4x + 6x - 24
:
y = x^2 + 2x - 24; is the equation